OpenStudy (anonymous):

What are the steps needed to solve this? sin^2 20degrees+1/sec^2 20degrees

OpenStudy (mathmale):

Unsure of what you mean by sin^20degrees; I think you mean sin (20 deg). Using that " ^ " here is very unusual. Oh. You meant (sin 20 deg)^2. Thanks for the clarification.

OpenStudy (mathmale):

\[\sin ^{2}x+\frac{ 1 }{ \sec ^{2}x }\] can be greatly simplified by using the common identity for \[\frac{ 1 }{ \sec x }\] . Try it. Write a new identity.